if the sides of a square are lengthned by 7 inches,the area becomes 121 in^2.find the length of a side of the original square.

Found 2 solutions by edjones, stanbon:
Answer by edjones(8007)

sqrt(121)=11
11-7=4" the length of a side of the original square.
.
Ed

Answer by stanbon(75887)

if the sides of a square are lengthned by 7 inches,the area becomes 121 in^2.find the length of a side of the original square.
-----
Area of a square = (side)^2
------------------
Equation:
(s+7)^2 = 121
s+7 = 11
side = 4 inches
----------------------
Cheers,
Stan H.
======================